How many combinations of 4 numbers can be made from the set of numbers {2,4,6,7,8,9}

a group consists of 6 men and 4 women. If a committee of 5 people is formed, find the number of combinations of the committee that has 3 men and 2 women.

Answer by stanbon(75887)

How many combinations of 4 numbers can be made from the set of numbers {2,4,6,7,8,9}
Ans 6C4 = 6C2 = (6*5)/(1*2) = 15 combinations
==================================================
a group consists of 6 men and 4 women. If a committee of 5 people is formed, find the number of combinations of the committee that has 3 men and 2 women.
---------------
Ans: 6C3 * 4C2 = 20 * 6 = 120
===================================
Cheers,
Stan H.
==================